OpenStudy (anonymous):

Help Solve the system by substitution. y=x-1 y=-2x+5 Could you show me how to these types of problems? Thanks

OpenStudy (opcode):

\[y=x-1\] \[y=-2x+5 \implies -2x+5 = x-1\] \[-2x+5 = x-1\] \[-x+5 = -1\] \[-x = -6\] \[x = 6\] Can you go from there?

OpenStudy (opcode):

Hint: You need to solve for \(y\).

OpenStudy (anonymous):

So from where you left it I would replace x for 6 in the original equation?

OpenStudy (opcode):

Oops, I made an error \(x = 2\) sorry, I am a bit sleeply. Yes, you now insert \(x\) into the original equation. \[y=2-1\] \[y=1\] \[1=-2(2)+5\] \[1=-4+5\] \[1=1\]

OpenStudy (anonymous):

I think I understand how to do these now. Thank you for the help.

OpenStudy (opcode):

Yeah, they tend to be easy, it's just late where I live so I made a mistake, sorry :-P.